OpenStudy (anonymous):

Solve the equation by finding the exact solution. ln x − ln 5 = 15

OpenStudy (anonymous):

\[x-5=e ^{15}\] \[x=e ^{15}+5\]

OpenStudy (ranga):

\[\Large \ln(A) - \ln(B) = \ln(\frac{ A }{ B })\]

OpenStudy (ranga):

\[\Large \ln(X) = C \quad means \quad X = e^c\]

OpenStudy (anonymous):

that's not my equation. my equation is this ln x − ln 5 = 15

OpenStudy (ranga):

I am giving you identities. Compare the first identity I gave you to the problem. Compare ln(A) - ln(B) to ln(x) - ln(5). What are A and B? Put that on the right hand side of the identity and tell me what you get.